An investment adviser invested $14,000 in two accounts. One investment earned 5% annual simple interest, and the other investment earned 2.5% annual simple interest. The amount of interest earned for 1 year was $535. How much was invested in each account?
Answer by ikleyn(52800)

.
An investment adviser invested $14,000 in two accounts.
One investment earned 5% annual simple interest,
and the other investment earned 2.5% annual simple interest.
The amount of interest earned for 1 year was $535.
How much was invested in each account?
~~~~~~~~~~~~~~~~~~~

Let x be amount invested at 5% annually.

Then the amount invested at 2.5% is the rest (14000-x) dollars.


5% investment generates the annual interest of 0.05x dollars.

2.5% investment generates the annual interest of 0.025*(14000-x).


The totall annual interest equation is

    0.05x + 0.025*(14000-x) = 535  dollars.


Simplify this equation and find x

    0.05x + 0.025*14000 - 0.025x = 535,

    0.05x - 0.025x = 535 - 0.025*14000

        0.025x     =     185

             x     =     185/0.025 = 7400.


Thus $7400 invested at 5%  and the rest,  $14000 - $7400 = $6600  invested at 0.025%.    ANSWER


CHECK.  0.05*7400 + 0.025*6600 = 535 dollars, the total annual interest.   ! correct !

Solved.

-----------------

To see other similar solved problems,  look into the lesson
    - Typical investment problems
    - Using systems of equations to solve problems on investment
in this site.